# Public-private funding models in open source software development: A case study on scikit-learn

### v4

@@ 4.3 Private Sponsorship: The Role of the scikit-learn Consortium
scikit-learn has received significant support from the private sector through the scikit-learn consortium and sponsorship of maintainers by companies like Nvidia, HuggingFace, and QuantSight. By creating the consortium, François Goupil explained they had sought to provide a more stable source of funding beyond research grants and to hire full-time maintainers, which was not possible previously.
@@
Companies in the consortium have pursued a range of social, economic, and technological goals. By joining the consortium, companies are able to support a software tool that they use, whilst building goodwill and a reputation as OSS advocates, which an important signal for hiring in a competitive labour market. François Goupil explained, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ’Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work”. He viewed this incentive as more important for companies that did not already have a strong reputation in ML/AI field, since it would make them more credible and attractive to job-seekers.
@@
For their part, consortium members cited a mix of pragmatic and principled reasons for funding scikit-learn. On a basic level, it allows them to support the maintenance of a tool they use in their products and services. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps [him] up at night”, citing numpy as a critical library that had been under-maintained for some time. Leo Dreyfus-Schmidt from Dataiku expressed a similar sentiment, who stated, “We are very much aware that we rely on open tools like scikit-learn” and highlighted the importance of “giv[ing] back to the community”. An anonymous consortium member was even more direct: “Companies pay millions in commercial licenses, so why not support the OSS projects we use¿‘
@@
However, discussions with maintainers and consortium members underlined that consortium membership is not simply a charitable contribution. In exchange for their sponsorship, companies gain a voice in scikit-learn’s development through bodies like the Technical Committee and Advisory Board, which are venues for surfacing industry use cases and needs. Leo Dreyfus-Schmidt characterised this as a “win-win”, allowing scikit-learn to benefit from “the perspective of people in industry, which…is a very different type of feedback from the one they get from the open source community”, whilst giving commercial members direct input into the project.
@@
In practice, this dynamic has yielded a mix of synergies and tensions. In some cases, companies’ business interests are compatible with scikit-learn’s technical roadmap. For example, in addition to its consortium membership, Nvidia has sponsored a maintainer to build a plug-in that enables alternative computational backends for scikit-learn. Mike McCarty explained, “One thing that Nvidia is very interested in is scikit-learn’s effort to accelerate with GPUs. I would love to see at the end of the day scikit-learn to be accelerated natively with a very nice user experience, kind of like what we have with PyTorch, which is the canonical example of a library that you can install and if you happen to have a GPU, you can accelerate it without changing your code”. Guillaume Lemaitre, a maintainer, described Nvidia’s simultaneous sponsorship of a maintainer and consortium membership as “a win-win situation”.
@@
In other cases, however, sponsor requests to shape the project’s direction or burnish their own reputations have run afoul of project norms. Two anonymous maintainers explained that in the past, a sponsor had requested modifications to the technical documentation that amounted, in their view, to free advertising for the company. This self-promotional request, they explained, was not acceptable, especially not in the technical documentation of the project. While disputes of this kind have been rare, one anonymous maintainer mentioned that disputes had led to the termination of sponsorship by one company in the past. To manage the interests and demands of consortium members, the scikit-learn maintainers employ carefully designed governance protocols that limit funders’ influence and prioritise community control. The Technical Committee, for instance, is barred from overriding decisions that achieve rough consensus amongst the maintainers. Adrin Jalali explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…because the power is given to the [maintainers] and that’s by design”.
@@
These governance measures are important to maintain the trust of the community and to mitigate fears that the project had been bought out. François Goupil explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop on MLOps because this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way…but it’s a discussion and in the end the team will decide if it’s worth it or not”, he explained. Through these governance protocols, the maintainers ensure that the project remains community-driven while still considering and incorporating feedback from the consortium. As François Goupil put it, “We are a community with different interests”, and preserving that pluralism is key to the project’s long-term sustainability and independence.
@@
Finally, even if relationships with their sponsors can be troublesome at times, the maintainers were grateful for their support and for not free-riding on scikit-learn, which many companies do. During an on-site visit, François Goupil showed me a dashboard, comprising world maps of contributors and charts of contributions from companies. Upon observing that the most active organisations were Chinese and American companies (see Table 2), I asked if these insights had influenced his opinion about whether they contributed sufficiently. He responded frankly: while such companies contribute useful issues and pull requests, their contributions ultimately create more work for the maintainers. He explained that it would be more helpful if these companies sponsored a maintainer or joined the consortium in order to provide more capacity to the maintainers to process the extra workload created by these contributions.
@@ 5.1 Implications for Research
5.1.1 The design and role of public-private OSS funding models
This study contributes novel findings on the roles of public and private funding in OSS development and maintenance. It reveals the diverse interests of funders, from the economic and technological goals of companies [48] to the policy objectives like digital sovereignty [1], national competitive in science and innovation [3], and economic growth [72] for governments. While the various stakeholders pursue different goals, they are not necessarily in conflict, as they share the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not necessarily share the government’s policy goals or digital sovereignty concerns, its long-term investment in scikit-learn aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government officials do not necessarily share the companies’ business interests in scikit-learn, the project benefits from companies’ feedback and funding for development and maintenance. The study also highlights the role of the governance mechanisms to balance the interests of the community and the companies in the consortium, extending prior work on the role of foundations as “boundary organisations” that mitigate the dominance of companies and enable collaboration between OSS developers and companies [16, 54].
5.1.2 Research design for future research on OSS funding
There is still a dearth of research on OSS funding, relative to other topics on the economics of OSS [91]. With the emergence of new funding models, from FOSS Contributor Funds to governmental funds like Germany’s STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s), amongst others. To this end, this paper contributes novel insights and a research design that can inform future studies on OSS funding. As discussed below, while the uniqueness of scikit-learn limits the generalisability of the case-specific findings, there are aspects of the scikit-learn experience that will resonate with OSS developers in general, such as the processes of balancing the varied interests of funders while preserving community values.
@@ 5.2 Implications for Practice
5.2.1 Implications for OSS Developer Communities
The case of scikit-learn shows that a diversified funding model can be an effective strategy for safeguarding a project’s independence and community-driven ethos whilst benefiting from both public and private support. Furthermore, public and private funding come with their own challenges. On the one hand, commercial sponsors often focus on specific technical issues related to maintenance and feature development, and their demands may sometimes conflict with project goals or norms. On the other hand, government funders may prioritise high-level policy objectives over the activities or expertise of the project. Navigating these tensions requires ongoing dialogue and negotiation to align goals and ensure a sustainable balance of funding sources. Ultimately, maintainers must be prepared to advocate for and defend their project’s needs and values. As the scikit-learn case demonstrates, a diverse funding portfolio can provide both stability and leverage in these discussions.
5.2.2 Implications for Companies
@@
5.2.3 Implications for Governments
For governments, the findings highlight the importance of funding the maintenance of existing, widely-used OSS infrastructure and not just the development of new libraries or features, extending prior recommendations for the funding of scientific OSS [68]. To ensure that funding is aligned with project needs, governments should engage in multi-stakeholder dialogue with OSS developers. This consultation is crucial for designing grants that balance policy goals with the expertise of maintainers. In the case of scikit-learn, initial government requests around expanding into deep learning were misaligned with the project’s scope and capabilities. Ongoing conversations were necessary to bridge this gap and ensure that funding supported both overarching policy objectives and essential project functions. Governments should proactively seek input from the OSS community to avoid such misalignments and to craft funding programs that effectively support the sustainability of critical (scientific) OSS projects.
@@ 5.3.1 Credibility
Credibility refers to the believability of qualitative research findings [93]. A number of steps were taken to enhance the credibility of the findings. First, multiple data sources—interviews, fieldwork, and secondary documents—were combined to triangulate findings and minimise the bias of individual sources [78]. Second, the prolonged engagement with the project’s maintainers over 17 months provided in-depth insights into the project’s evolving funding model, witnessing the ongoing negotiation with the government, as well as the arrival and departure of different consortium members. This longitudinal approach minimised the biases of snapshots at one moment in time. Third, findings were member-checked with research participants to enhance their accuracy [88]. Fourth, the community manager was invited to review a draft of the case study, further enhancing the accuracy and resonance of the findings [78]. However, data collection was constrained by data access challenges. A major obstacle was gaining access to interviewees, especially high-ranking stakeholders in the government and at sponsor companies, which is a common problem in “elite interviewing” [94]. For example, despite the generous support of the community manager and the offer of gift vouchers, only three representatives of commercial sponsors could be recruited for interviews.
@@ 5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefiting from the feedback from domain experts as well as peer review. The combination of these steps enhanced the dependability of the study.
@@ 6 Conclusion
This study on the funding model of scikit-learn makes two key contributions to research and practice. First, it contributes novel findings on emerging OSS funding models, illustrating how the maintainers of an impactful community-led OSS project have combined public and private funding to sustain the development and maintenance of the project, whilst safeguarding its community ethos. Second, it offers practical lessons for the funding of community-led OSS projects, and it makes recommendations to various OSS practitioners. In particular, it highlights effectiveness of diversified funding models as a method to sustain community-led OSS projects in industry-dominated R&D fields, such as AI.

### v5

@@ 4.3 Private Sponsorship: The Role of the scikit-learn Consortium
Companies in the consortium have pursued a range of social, economic, and technological goals. Goupil explained that companies typically join the consortium to support the maintenance of scikit-learn which they use in their products or services, as well as to build goodwill in the community, which can be useful for hiring ML researchers and developers. He commented, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ‘Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work.”’ He suggested this incentive may be more important for companies that do not already have a strong reputation in ML or AI, making them more credible and attractive to job-seekers.
@@
For their part, consortium members cited a mix of pragmatic and principled reasons for funding scikit-learn. On a basic level, it allows them to support the maintenance of a tool they use and rely on. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps [him] up at night,” citing numpy as a critical library that had been under-maintained for some time. Leo Dreyfus-Schmidt from Dataiku expressed a similar sentiment when he stated that, “We are very much aware that we rely on open tools like scikit-learn” and highlighted the importance of “giv[ing] back to the community.” An anonymous consortium member was even more direct: “Companies pay millions in commercial licenses, so why not support the OSS projects we use?”
@@
Of course, consortium membership is not simply a charitable contribution. In exchange for their sponsorship, companies gain a voice in scikit-learn’s development through bodies like the Technical Committee and Advisory Board, which are venues for surfacing industry use cases and needs. Dreyfus-Schmidt characterised the interactions in these committees as a “win-win” for the maintainers and the companies, allowing scikit-learn to benefit from “the perspective of people in industry, which…is a very different type of feedback from the one they get from the open source community,” whilst giving commercial members direct input into the project.
@@
In practice, this dynamic has yielded synergies and tensions. Largely, companies’ interests have been compatible with those of the maintainers. For example, Nvidia has sponsored a maintainer to build a plug-in enabling alternative computational backends for scikit-learn. McCarty explained, “Nvidia is very interested in scikit-learn’s effort to accelerate with GPUs. I would love to see scikit-learn accelerated natively with a nice user experience, like PyTorch, which you can install and accelerate without changing code.” Guillaume Lemaitre, a maintainer, described Nvidia’s simultaneous sponsorship and consortium membership as “a win-win situation” for the maintainers. However, in other cases, sponsors’ requests to shape the project’s direction or to burnish their reputations have run afoul of project norms. Two maintainers explained that a sponsor had once requested modifications to the technical documentation, which amounted in their view to free advertising of their products, which was unacceptable to the maintainers. While disputes of this kind have been rare, a dispute had led to the termination of a consortium membership in the past.
@@
To balance the diverse interests of their community and the member companies in the consortium, the scikit-learn maintainers employ governance protocols that limit funders’ influence and prioritise community control. The Technical Committee, for instance, is barred from overriding decisions that achieve rough consensus amongst the maintainers. Jalali explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…because the power is given to the [maintainers] and that’s by design.”
@@
These governance measures are important to maintain the trust of the community and to mitigate fears that the project has been bought out. Goupil explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop on MLOps because this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way…but it’s a discussion and in the end the team will decide if it’s worth it or not,” he explained. Through these governance protocols, the maintainers ensure that the project remains community-driven while still considering and incorporating feedback from the consortium. As Goupil put it, “We are a community with different interests,” and preserving that pluralism is key to the project’s long-term sustainability and independence.
@@
Finally, even though relationships with companies can be troublesome at times, the maintainers were generally grateful for their support and for not simply free-riding on scikit-learn, which many companies do. For example, during an on-site visit, Goupil showed me a dashboard, comprising charts of contributions from companies. Upon observing that the most active organisations are Chinese and American companies (see Table 2), I asked if these insights had influenced his opinion about whether they contribute sufficiently to the project. He responded frankly: while such companies contribute useful issues and pull requests, at the end of the day their contributions create more work for the maintainers. He explained that it would be more helpful if these companies sponsored a maintainer or joined the consortium in order to provide more capacity to the maintainers to process the extra workload created by these contributions.
@@ 5.1 Implications for Research
5.1.1 Relative Benefits & Drawbacks of OSS Funding Sources
This study contributes novel findings on how OSS developers view the relative benefits and drawbacks of public and private funding. In the case of scikit-learn, private funding has been associated with an understanding of the importance of maintenance and it has introduced the maintainers to industry perspectives and use cases, but it has also required careful management of sponsors’ interests. By contrast, the sizeable government grant has enabled long-term planning around hiring and the technical roadmap, but it has posed challenges in aligning policy goals with OSS realities, balancing national interests with community norms, and navigating the slow pace of disbursement. Overall, this case study shows that a diversified funding model can be an effective strategy for balancing the relative benefits and drawbacks of public and private funding. As Goupil explained, “Overall, it’s good to have a panaché.”
The study also contributes novel insights on the diverse interests of OSS funders, from the economic and technological goals of companies [48] to the policy objectives of governments like digital sovereignty [1], national competitive in science and innovation [3], and economic growth [72]. While public and private funders pursue different goals, the case study illustrates that they are not necessarily in conflict, sharing the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not necessarily share the government’s policy goals, such as digital sovereignty, its long-term investment in scikit-learn aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government officials do not necessarily share the companies’ business interests in scikit-learn, the project benefits from companies’ feedback and funding for feature development and maintenance.
5.1.2 Designing Future Research on OSS Funding
There is a dearth of research on OSS funding, relative to other topics on the economics of OSS [91]. With new funding models emerging, from FOSS Contributor Funds to governmental funds like Germany’s STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s). Future research should expand this study by investigating the design and outcomes of different governmental OSS funding models, such as Germany’s STF or the European Commission’s Next Generation Internet initiative. By examining these aspects, researchers can contribute to the development of best practices and guidelines for the design and implementation of governmental funding programmes that effectively support the sustainability of OSS projects.
@@ 5.2 Implications for Practice
5.2.1 Recommendations for OSS Developers
The case of scikit-learn shows that a diversified funding model can be an effective strategy for safeguarding a project’s independence whilst benefiting from both public and private support. Furthermore, public and private funding come with their unique benefits and drawbacks. On the one hand, private funders understand the importance of maintenance and share industry use cases that lead to the development of new features. However, their demands can conflict with project goals or norms. On the other hand, public funding can provide sizeable grants, which support long-term planning but may prioritise innovation over maintenance or policy goals over the specialism or needs of the project. Maintainers must be prepared to advocate for and defend their project’s needs and community values, which are aided by project governance protocols and, if possible, regular dialogue during the design of grants. A diverse funding portfolio can provide both stability and leverage in these discussions.
5.2.2 Recommendations for Companies
@@
5.2.3 Recommendations for Governments
For governments, the findings extend prior recommendations for the funding of scientific OSS [68] by highlighting the importance of funding the maintenance of existing OSS in addition to new innovation. To ensure that funding is aligned with project needs, governments should engage in multi-stakeholder dialogue with OSS developers. This consultation is crucial for designing grants that balance policy goals with the expertise of maintainers. In the case of scikit-learn, initial government requests around expanding into deep learning were misaligned with the project’s scope and capabilities. Ongoing conversations were necessary to bridge this gap and ensure that funding supported both overarching policy objectives and essential project functions. Governments should proactively seek input from the OSS community to avoid misalignments and to craft funding programs that effectively support the needs of critical OSS projects.
@@ 5.3.1 Credibility
Credibility refers to the believability of qualitative research findings [93]. Multiple data sources—interviews, fieldwork, and secondary documents—were combined to triangulate findings and minimise bias [78]. The prolonged engagement with the project’s maintainers over 17 months provided in-depth insights, minimising the biases of snapshots at one moment in time. Findings were member-checked with participants and reviewed by the community manager to enhance accuracy [88, 78]. However, data collection was constrained by access challenges, especially with high-ranking stakeholders in the government and sponsor companies, which is a common problem in “elite interviewing” [94]. For example, despite the outreach support of the community manager, only three representatives of commercial sponsors could be recruited for interviews. Despite these limitations, the triangulation of multiple data sources, prolonged engagement, and member checking helped to ensure the credibility and trustworthiness of the findings.
@@ 5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefiting from the feedback from domain experts as well as peer review.
@@ 6 Conclusion
This study on the funding model of scikit-learn makes two key contributions to research and practice. First, it contributes novel findings on emerging OSS funding models, illustrating how the maintainers of an impactful community-led OSS project have integrated and balanced public and private funding to sustain the development and maintenance of the project. Second, it offers practical lessons for the funding of community-led OSS projects, and it makes recommendations to various practitioners. In particular, it provides insights into the benefits and drawbacks of public and private funding for OSS, and how these different funding sources can be balanced to support project goals. Importantly, it highlights effectiveness of diversified funding models as a method to sustain community-led OSS projects in industry-dominated R&D fields like AI. Overall, the findings contribute to the literature on OSS funding and provide actionable recommendations for governments and companies to support the sustainability of OSS.